OpenStudy (itrymath):

The function f(x) is shown below. f(x) = (3)x - 2 The function f(x) is shifted to the left by 13 units. Which of the following best represents the new function? (4 points) f(x) = (3)x + 11 f(x) = (3)x + 13 f(x) = (3)x -15 f(x) = (3)x - 13

OpenStudy (itrymath):

@YanaSidlinskiy

OpenStudy (yanasidlinskiy):

HINT: to shift f(x) to the left by 13 units, you have to replace x with x-13.Evaluate: f(x-13)

OpenStudy (yanasidlinskiy):

Did you get your answer:)

OpenStudy (itrymath):

C?

OpenStudy (yanasidlinskiy):

I kinda just gave away the answer if you read closely.

OpenStudy (yanasidlinskiy):

"X-13"

OpenStudy (itrymath):

yeah but there is a 2...

OpenStudy (itrymath):

x-2 +13 = x-15

OpenStudy (yanasidlinskiy):

It's not asking to solve or is it?

OpenStudy (itrymath):

no

OpenStudy (itrymath):

but if we add 13 to the original function that has x-2 wouldnt we get x-15?

OpenStudy (itrymath):

OH im sos sorry !!!! you cant see it

OpenStudy (itrymath):

@YanaSidlinskiy you still there?

OpenStudy (itrymath):

@jim_thompson5910

OpenStudy (itrymath):

@ShadowLegendX

OpenStudy (itrymath):

isnt C right @ShadowLegendX

jimthompson5910 (jim_thompson5910):

to shift left 13 units, replace every x with (x+13) \[\Large f(x) = (3)^{x-2}\] \[\Large f(x+13) = (3)^{x+13-2}\] \[\Large f(x+13) = (3)^{x+11}\]

OpenStudy (itrymath):

OOOH i see

OpenStudy (itrymath):

omg lol so simple and i didnt know xD thanks @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no problem